The equation of the line that goes through the point(3,4) and is parallel to the line going through the points (-6,6) and (1,3) can be written in the form y=mx+b where
m=
b=
I am having difficulty with the second part only m=-3/7

Answer by nerdybill(5411)

The equation of the line that goes through the point(3,4) and is parallel to the line going through the points (-6,6) and (1,3) can be written in the form y=mx+b where
m=
b=
.
You already discover that the slope of a line going through the points (-6,6) and (1,3) is
m = -3/7
.
Now, you have a "point" (3,4) and a "slope" (-3/7).
Plug it all into the "point-slope" form of a line:
y – y1 = m(x – x1)
.
y – 4 = (-3/7)(x – 3)
7y – 28 = (-3)(x – 3)
7y – 28 = -3x + 9
7y = -3x + 37
.
Your answer then is:
y = (-3/7)x + (37/7)
m= -3/7
b= 37/7